OpenStudy (anonymous):

What are the factors of 3x^2 + 13x + 4

OpenStudy (anonymous):

Polynomials are factored by their roots. That is, when f(x) = 0, the values of x can be used to factor the polynomial.

OpenStudy (anonymous):

To find the root of a quadratic equation in the form of \(ax^2+bx+c\), it's easiest to use the quadratic equation to find the roots.

OpenStudy (anonymous):

Once you have found the roots, the polynomial is factored as \((x-r_1)(x-r_2)=0\).

OpenStudy (anonymous):

So if the roots are 1 and -2, then it's factored as \((x-1)(x+2)\)

OpenStudy (anonymous):

????? i dont get it

OpenStudy (anonymous):

there are multiple ways to factor a polynomial

OpenStudy (anonymous):

what method are you supposed to use?

OpenStudy (anonymous):

Hint: 3*4=12 1*12=12 1+12=13.

OpenStudy (anonymous):

Do you know how to do factor-by-grouping? \( 3x^2 + 13x + 4 = 3x^2+12x+x+4\)

OpenStudy (anonymous):

no i dont understand it

OpenStudy (anonymous):

Factoring a quadratic trinomial is the reverse of multiplying linear binomials using the FOIL method (First Outer Inner Last)

OpenStudy (anonymous):

If you multiply something like (x+3(2x-2), you get 2x^2 -2x +6x -6. After combining like terms, you'll have the trinomial, 2x^2+4x-6. You can see that if you do it in reverse, you can find two numbers that multiply to make 12 (2*6), that add to make 4.

OpenStudy (anonymous):

(1x+4)(3x+1) =3x^2+1x+12x+4= 3x^2+13x+4